OpenStudy (anonymous):

What is the area ABCD with vertices A(-4,-6) B(6, -6) C(-1, 5) D(9, 5)

OpenStudy (anonymous):

okay. So you can either graph or use the distance formula and you have to find the distance between A and B and B and C then multiply. DO you know how to do that?

OpenStudy (anonymous):

whats the formula?

OpenStudy (anonymous):

@ab60093

OpenStudy (anonymous):

oh. Well, personally its easier to graph but the distance formula is \[d=\sqrt{(x1-x2)^{2}+(y1-y2)^{2}}\]

OpenStudy (anonymous):

so i do that to each one? will it be the final answer?

OpenStudy (anonymous):

You would only have to do it to AB and BC and no that wouldn't be the answer.so it would be set up like this for AB \[d=\sqrt{(-4-6)^{2}+(-6-(-6))^{2}}\]

OpenStudy (anonymous):

and BC would be something like this \[BC=\sqrt{(6-(-1))^{2}+(-6-5)^{2}}\]

OpenStudy (anonymous):

for AB its \[AB=\sqrt{(10)^{2}+(0)^{2}}\]

OpenStudy (anonymous):

so the answer would be square root of 100?

OpenStudy (anonymous):

for AB

OpenStudy (anonymous):

yes. For AB

OpenStudy (anonymous):

okay now for CD i do the same like AB?

OpenStudy (anonymous):

Not quite. You would do BC not CD. And i wrote the equation up top

OpenStudy (anonymous):

so it would be (7)^2+(11)^2?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

so its \[\sqrt{170}\]

OpenStudy (anonymous):

i got 178

OpenStudy (anonymous):

its 7x7 which is 49 plus 11x11 which is 121. 121+49=170

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

am i done yet?

OpenStudy (anonymous):

not quite. You must now multiply the lengths of AB and BC. so 10 x square root of 170

OpenStudy (anonymous):

THAT would be the answer

OpenStudy (anonymous):

130.384048104?

OpenStudy (anonymous):

yeah. or just \[10\sqrt{170}\]

OpenStudy (anonymous):

another way to do this would have been the Pythagorean theorem but that would have been the same amount of work

OpenStudy (anonymous):

omg thank you so much!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

i have one more problem though

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what is the area of DEF with vertices D(-1, -5) E(4, -5) F(4, 7)

OpenStudy (anonymous):

ooo. Triangle. fun

OpenStudy (anonymous):

ahaha

OpenStudy (anonymous):

okay. this one is going to be a little bit different to make our lives easier. We are going to use the graphing style. so when you do this you get a right triangle. do you know how to find the area of a right triangle?

OpenStudy (anonymous):

A=ab 2?

OpenStudy (anonymous):

a=ab/2?

OpenStudy (anonymous):

yep. hold on like 2 seconds though

OpenStudy (anonymous):

okie

OpenStudy (anonymous):

okay. Sorry i was ordering food. ANYWAYS, uhhh the distance between E and F is 12 because...it just is. When you look at a graph the distance is 12, if you want we can do the distance formula to check later. The distance between D and E is 5. You would now have to set up the formula \[A=\frac{ d \times e }{ 2 }\]

OpenStudy (anonymous):

wait oops

OpenStudy (anonymous):

\[A=\frac{ DE \times EF }{ 2 }\]

OpenStudy (anonymous):

so input values and solve

OpenStudy (anonymous):

okay hold on let me solve this lol

OpenStudy (anonymous):

so for D do i put -1 or -5?

OpenStudy (anonymous):

@ab60093

OpenStudy (anonymous):

oh, lol sorry

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

it would be \[A=\frac{ (5\times 12) }{ 2 }\]

OpenStudy (anonymous):

You wouldn't use the coordinates in the equation. You would use the distance between the coordinates

OpenStudy (anonymous):

30?

OpenStudy (anonymous):